payton:

A pair of linear equations is shown: y = −x + 1 y = 2x + 4 Which of the following statements best explains the steps to solve the pair of equations graphically? A. On a graph, plot the line y = −x + 1, which has y-intercept = −1 and slope = 1, and y = 2x + 4, which has y-intercept = 2 and slope = 4, and write the coordinates of the point of intersection of the two lines as the solution. B. On a graph, plot the line y = −x + 1, which has y-intercept = 1 and slope = 1, and y = 2x + 4, which has y-intercept = 1 and slope = 4, and write the coordinates of the point of intersection of the two lines as the solution. C. On a graph, plot the line y = −x + 1, which has y-intercept = 1 and slope = −1, and y = 2x + 4, which has y-intercept = −2 and slope = 2, and write the coordinates of the point of intersection of the two lines as the solution. D. On a graph, plot the line y = −x + 1, which has y-intercept = 1 and slope = −1, and y = 2x + 4, which has y-intercept = 4 and slope = 2, and write the coordinates of the point of intersection of the two lines as the solution.

ramen:

A nd B are wrong cuz the slope or intercepts are stated wrong

jimthompson5910:

y = mx+b m = slope b = y intercept For example, y = 7x+12 has slope m = 7 and y intercept b = 12

payton:

okay

payton:

C?

ramen:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jimthompson5910 y = mx+b m = slope b = y intercept For example, y = 7x+12 has slope m = 7 and y intercept b = 12 \(\color{#0cbb34}{\text{End of Quote}}\) Based on this payt, which one has stated the correct slope and intercepts?

ramen:

For c, is the y intercept -2?

payton:

yes

payton:

¨which has y-intercept = −2 and slope = 2¨

ramen:

For the equation y=2x+4, is the y intercept -2?

payton:

oh no

payton:

so c isnt my answer then

payton:

it has to be D

ramen:

Yes correct